OpenStudy (love.kat123):

How much has the value of y changed between the points (–8, 4) and (8, –4)? A. 1/2 units B. 2 units C. 8 units D. 16 units

OpenStudy (love.kat123):

@Adjax

OpenStudy (anonymous):

see the slope = \[\frac{ \Delta y }{ \Delta x }\] Work this out on this formula

OpenStudy (love.kat123):

okay \[\frac{ -8,4}{ 8,-4 }\]

OpenStudy (anonymous):

nah you haven't got that right \[\frac{ \Delta y }{ \Delta x } = y_{2} - y_{1} /x_{2} - x_{1}\] Here P1 = (x1,y1) = (-8,4)( P2 = (x2,y2) = (8,-4) Therefore: \[\frac{ \Delta y }{ \Delta x } =\frac{ 4 - (-4) }{ 8 - (-8)} =\frac{ 8 }{16} = \frac{ 1 }{ 2 }\] So the answer is 1/2 unit change @Love.Kat123

OpenStudy (love.kat123):

oh :/ i suck at this thanks